OpenStudy (anonymous):

find the integral : cos3xcosx dx thanks.

OpenStudy (anonymous):

Use the following identities: \[\cos(a+b)x=\cos ax\cos bx-\sin ax \sin bx\\ \cos(a-b)x=\cos ax\cos bx+\sin ax\sin bx\] Add the two together, you have \[\cos(a+b)x+\cos(a-b)x=2\cos ax\cos bx\\ \cos ax\cos bx=\frac{1}{2}(\cos(a+b)x+\cos(a-b)x)\] where \(a=3\) and \(b=1\).